# Composition of Scalar Mesons and Their Effects on Nuclear Matter Properties in an Extended Linear Sigma Model

### v2

@@ III Conclusion and discussion
In this work, we continued to study NM properties and NS structures by extending the bELSM previousely proposed in Ref. Ma and Ma (2024) to include a new type of flavor structure which is constructed with an antisymmetric tensor of the corresponding baryon and meson representations. With this extension, the bELSM at the lowest order could reproduce the empirical spectra of the scalar mesons and nucleons, as well as the NM properties at saturation density.
@@
Explicitly, we found that without introducing the 12-quark terms as in the previous work, the reasonable behavior of nuclear matter at high densities can be well reproduced. The couplings related to $\rho$ and $\omega$ mesons can be separated due to the new flavor structures, allowing the $E_{\rm sym}$ plateau to appear at the expected density range, around $2n_{0}$, instead of below $n_{0}$ as in the previous work. This is consistent with Ref. Li et al. (2022), which highlights the importance of such $E_{\rm sym}$ behavior to meet the constraints of GW170817 and the neutron skin thickness of ${}^{208}\rm Pb$ with a unified model. Besides the phenomena at densities, the pion decay constant is also consistent with recent lattice QCD results Aoki et al. (2020), with $f_{\pi}\approx 134~{}\rm MeV$ in el-g30eg, compared to $f_{\pi}\sim(60-70)~{}\rm MeV$ in the previous work.
@@
With this improved lowest-order bELSM and the corresponding parameter space, a further study on the EOS over a wider density region was performed. Several parameter sets were found to meet the NM properties at $n_{0}$. The differences lie in the magnitude of $g_{aNN}$ and the four-vector meson coupling, $\tilde{g}_{3}$. It was found that small $g_{aNN}$ and $\tilde{g}_{3}$ lead to more compact NSs, consistent with the constraints of MSP J0740+6620 and GW170817. The influences of the EOS behaviors on NS M-R relations were also discussed: a soft $E_{\rm sym}$ at intermediate densities, with a stiff behavior at subsaturation densities, is necessary for the compactness of NSs to meet the GW170817 constraints; a stiff SV behavior at high densities is necessary for the $M_{\rm max}$ of NSs to meet the MSP J0740+6620 constraints. Besides the well-produced NS structures from el-g30eg and FSU-$\delta 6.7$, the SV behavior of el-g30eg is more favored, as its SV behavior can approach the conformal limit at high densities, expected for the fact that QCD is asymptoticly free at high energies.
@@
The percentages of 2-quark and 4-quark constituents making up the scalar mesons were also calculated. The $\sigma$ meson was found to be composed of 2-quark and 4-quark constituents in similar proportions, while the $a_{0}$ meson is mainly composed of 4-quark constituents in the current optimal parameter spaces. This is consistent with previous analyses in vacuum Fariborz et al. (2008). The variation of the constituents due to different magnitudes of $g_{aNN}$ and four-vector meson couplings were also provided for future studies on scalar mesons and the macroscopic properties of compact objects in the universe. However, the quark constituents make the couplings related to $a_{0}$, both types: $g_{a_{0}NN}$ and $C_{\sigma^{i}\delta^{2}}$, different from the results of FSU-$\delta 6.7$ and Zabari-19, which are used for the symmetry energy at high densities. This again signifies the importance of considering the chiral symmetry pattern of QCD at low energies in phenomenological analyses beyond the saturation density.
@@
In summary, the lowest-lying isoscalar and isovector scalar and vector mesons are all included within a unified model, bELSM, which also respects the chiral symmetry pattern of QCD at low energies. The NM properties and NS structures are systematically studied. Our recent works Guo et al. (2024a); Ma and Ma (2024); Guo et al. (2024b); Zhang et al. (2024a, b) focus on the relationships between the QCD symmetry pattern and compact objects in the universe. In the near future, improved AI analysis on a general Walecka-type model with the inclusion of all types of lowest-lying mesons and baryons will be available, which will help further investigate bELSM at densities due to its numerical complexity. Additionally, studies based on bELSM with the inclusion of strange freedoms are underway to address the properties of compact objects with hyperons. Meanwhile, the Hartree-Fock approach with $\chi\rm EFT$ will also be initially completed to study the EOS of NM, including quantum effects, which are ignored in the current RMF approximation.
@@ Acknowledgements.
Appendix A The lowest order bELSM
In this appendix, we provide details of the bELSM at the lowest order before the RMF approximation. In the model, the scalar mesons $S$ and pseudoscalar mesons $P$ are combined into a compact form $\Phi=S+iP$:
@@ Appendix A The lowest order bELSM
The representations of the 2-quark and 4-quark (pseudo-)scalar configurations are similar, but with ”prime” and ”hat-prime” indices, as defined in the main text. For the vector and axial-vector mesons, they are expressed as:
@@
where $L(R)^{\mu}$ stands for the left(right)-handed vector mesons, and $V(A)^{\mu}$ denotes the vector(axial-vector) mesons. The baryon octet has expression:
@@
The chiral symmetry transformations for (pseudo-)scalar configurations are given by Fariborz et al. (2005, 2008, 2009):
@@
where $g_{\mathrm{L},\mathrm{R}}\in\mathrm{SU}(3)_{\mathrm{L},\mathrm{R}}$. They remain unchanged for $\rm U(1)_{V}$ transformation. But for $\rm U(1)_{A}$ transformation, they transform as:
@@
with $\nu$ being the chiral angle, and $q\bar{q}$ and $(qq)(\bar{q}\bar{q})$ configurations are assumed for 2-quark and 4-quark components, respectively. For the vector and axial-vector mesons, their chiral transformations are Parganlija et al. (2013):
@@
In this work, the diquark approximation for baryons is adopted Olbrich (2015); Olbrich et al. (2016). The baryons are denoted as $N_{\mathrm{R}}^{(\mathrm{RR})}$, $N_{\mathrm{L}}^{(\mathrm{LL})}$, $N_{\mathrm{L}}^{(\mathrm{RR})}$, and $N_{\mathrm{R}}^{(\mathrm{LL})}$, where the superscripts represent the quark configuration and the subscripts represent the diquark configuration. The chiral transformations for baryons are:
@@
Before constructing the full Lagrangian, a series of power-counting rules are needed to organize the effective operators:
@@
Consequently, the full Lagrangian at the lowest order is
@@
The (pseudo-)scalar meson part is Fariborz et al. (2008):
@@
$\mathcal{L}_{\rm B}$ $=$ ${\rm Tr}\left(\bar{B}i\not{\partial}B\right)+c{\rm Tr}\left(\bar{B}\gamma_{\mu}V^{\mu}B+\bar{B}\gamma_{\mu}\gamma_{5}A^{\mu}B\right)+c^{\prime}{\rm Tr}\left(\bar{B}\gamma_{\mu}BV^{\mu}+\bar{B}\gamma_{\mu}\gamma_{5}BA^{\mu}\right)$ (24)
@@ Appendix B Approximations in the current work
In the present study of NM properties, we take RMF, which is a classical approximation, where meson fields are treated as potentials induced by matter fields, such as nucleons in the current work. Additionally, the system composed of these baryons is considered to be infinite, homogeneous, and static, so that any quantum numbers and currents are averaged to zero. This results in the following approximations
@@
and other related meson fields are also treated similarly. For instance, $K$ mesons are excluded due to their non-zero strange quantum number.
Mesons heavier than $1\rm GeV$, such as $a^{\prime}_{0}$ and $\omega_{S}$, are ignored since their contributions are suppressed at the current working scale. This leads to the following parameter choices in the current work: $\tilde{g}_{3}=2(g_{3}+g_{4})$, $\tilde{h}_{2}=2h_{1}+h_{2}$, while the other parameters remain the same as the original ones after the approximations. Details on how to solve the RMF equations of motion (EOMs) to obtain the equation of state (EOS) after these approximations are provided in our other work Guo et al. (2024b).

### v3

@@ III Conclusion
In this work, we continued to study NM properties and NS structures by extending the bELSM previousely proposed in Ref. Ma and Ma (2024) to include a new type of flavor structure which is constructed with an antisymmetric tensor of the corresponding baryon and meson representations. With this extension, the bELSM at the lowest-order could reproduce the empirical spectra of the scalar mesons and nucleons as well as the NM properties at saturation density.
@@
Explicitly, we found that without introducing the two-quark terms as in the previous work, the reasonable behavior of nuclear matter at high densities can be well reproduced. The couplings related to $\rho$ and $\omega$ mesons can be separated due to the new flavor structures, allowing the $E_{\rm sym}$ plateau to appear at the expected density range, around $2n_{0}$, instead of below $n_{0}$ as in the previous work. This is consistent with Ref. Li et al. (2022), which highlights the importance of such $E_{\rm sym}$ behavior to meet the constraints of GW170817 and the neutron skin thickness of $^{208}\rm Pb$ with a unified model. Besides the phenomena at densities, the pion decay constant is also consistent with recent lattice QCD results Aoki et al. (2020), with $f_{\pi}\approx 134~\rm MeV$ in el-g30eg, compared to $f_{\pi}\sim(60-70)~\rm MeV$ in the previous work.
@@
With this improved lowest-order bELSM and the corresponding parameter space, a further study on the EOS over a wider density region was performed. Several parameter sets were found to meet the NM properties at $n_{0}$. The differences lie in the magnitude of $g_{{a_{0}NN}}$ and the four-vector meson coupling, $\tilde{g}_{3}$. It was found that small $g_{{a_{0}NN}}$ and $\tilde{g}_{3}$ lead to more compact NSs, consistent with the constraints of MSP J0740+6620 and GW170817. The influences of the EOS behaviors on NS M-R relations were also discussed: a soft $E_{\rm sym}$ at intermediate densities, with a stiff behavior at subsaturation densities, is necessary for the compactness of NSs to meet the GW170817 constraints, and a stiff SV behavior at high densities is necessary for the $M_{\rm max}$ of NSs to meet the MSP J0740+6620 constraints. Besides the well-produced NS structures from el-g30eg and FSU-$\delta 6.7$, the SV behavior of el-g30eg is more favored, as its SV behavior can approach the conformal limit at high densities, expected for the fact that QCD is asymptotically free at high energies.
@@
The percentages of two-quark and four-quark constituents making up the scalar mesons were also calculated. The $\sigma$ meson was found to be composed of two-quark and four-quark constituents in similar proportions, while the $a_{0}$ meson is mainly composed of four-quark constituents in the current optimal parameter spaces. This is consistent with previous analyses in vacuum Fariborz et al. (2008). The variation of the constituents due to different magnitudes of $g_{{a_{0}NN}}$ and four-vector meson couplings were also provided for future studies on scalar mesons and the macroscopic properties of compact objects in the universe. However, the quark constituents make the couplings related to $a_{0}$, both types: $g_{{a_{0}NN}}$ and $C_{\sigma^{i}{a_{0}}^{2}}$, with i=1, 2i=1,\ 2, different from the results of FSU-$\delta 6.7$ and Zabari-19, which are used for the symmetry energy at high densities. This again signifies the importance of considering the chiral symmetry pattern of QCD at low energies in phenomenological analyses beyond the saturation density.
@@
In summary, the lowest-lying isoscalar and isovector scalar and vector mesons are all included within a unified model, bELSM, which also respects the chiral symmetry pattern of QCD at low energies. The NM properties and NS structures are systematically studied. Our recent works Guo et al. (2024a); Ma and Ma (2024); Guo et al. (2024b); Zhang et al. (2024a, b) focus on the relationships between the QCD symmetry pattern and compact objects in the universe. In the near future, improved AI analysis on a general Walecka-type model with the inclusion of all types of lowest-lying mesons and baryons will be available, which will help further investigate bELSM at finite densities due to its numerical complexity. Additionally, studies based on bELSM with the inclusion of strange freedoms are underway to address the properties of compact objects with hyperons. Meanwhile, the Hartree-Fock approach with chiral effective field theories will also be initially completed to study the EOS of NM, including quantum effects, which are ignored in the current RMF approximation.
@@ Acknowledgements.
Appendix A The lowest-order bELSM
In this Appendix, we provide details of the bELSM at the lowest-order before the RMF approximation. In the model, the scalar mesons $S$ and pseudoscalar mesons $P$ are combined into a compact form $\Phi=S+iP$:
@@ Appendix A The lowest-order bELSM
The representations of the two-quark and four-quark (pseudo-)scalar configurations are similar, but with ”prime” and ”hat-prime” indices, as defined in the main text. For the vector and axial-vector mesons, they are expressed as,
@@
where $L(R)^{\mu}$ stands for the left-(right)-handed vector mesons, and $V(A)^{\mu}$ denotes the vector(axial-vector) mesons. The baryon octet has expression:
@@
The chiral symmetry transformations for (pseudo)scalar configurations are given by Fariborz et al. (2005, 2008, 2009):
@@
where $g_{\mathrm{L},\mathrm{R}}\in\mathrm{SU}(3)_{\mathrm{L},\mathrm{R}}$. They remain unchanged for $\rm U(1)_{V}$ transformation. But for $\rm U(1)_{A}$ transformation, they transform as,
@@
with $\nu$ being the chiral angle, and $q\bar{q}$ and $(qq)(\bar{q}\bar{q})$ configurations are assumed for two-quark and four-quark components, respectively. For the vector and axial-vector mesons, their chiral transformations are Parganlija et al. (2013),
@@
In this work, the diquark approximation for baryons is adopted Olbrich (2015); Olbrich et al. (2016). The baryons are denoted as $N_{\mathrm{R}}^{(\mathrm{RR})}$, $N_{\mathrm{L}}^{(\mathrm{LL})}$, $N_{\mathrm{L}}^{(\mathrm{RR})}$, and $N_{\mathrm{R}}^{(\mathrm{LL})}$, where the superscripts represent the quark configuration and the subscripts represent the diquark configuration. The chiral transformations for baryons are
@@
Before constructing the full Lagrangian, a series of power counting rules is needed to organize the effective operators:
@@
Consequently, the full Lagrangian at the lowest-order is
@@
The (pseudo)scalar meson part is Fariborz et al. (2008)
@@
$\mathcal{L}_{\rm B}$ $=$ ${\rm Tr}\left(\bar{B}i\not{\partial}B\right)+c{\rm Tr}\left(\bar{B}\gamma_{\mu}V^{\mu}B+\bar{B}\gamma_{\mu}\gamma_{5}A^{\mu}B\right)+{c^{\prime}{\rm Tr}\left(\bar{B}\gamma_{\mu}BV^{\mu}\right)+c^{\prime}_{A}\left(\bar{B}\gamma_{\mu}\gamma_{5}BA^{\mu}\right)}$ (24)
@@ Appendix B Approximations in the current work
In the present study of NM properties, we take RMF, which is a classical approximation, where meson fields are treated as potentials induced by matter fields, such as nucleons in the current work. Additionally, the system composed of these baryons is considered to be infinite, homogeneous, and static, so that any quantum numbers and currents are averaged to zero. This results in the approximations
@@
and other related meson fields are also treated similarly. For instance, $K$ mesons are excluded due to their nonzero strange quantum number.
Mesons heavier than $1~\rm GeV$, such as $a^{\prime}_{0}$ and $\omega_{S}$, are ignored since their contributions are suppressed at the current working scale. This leads to the following parameter choices in the current work: $\tilde{g}_{3}=2(g_{3}+g_{4})$, $\tilde{h}_{2}=2h_{1}+h_{2}$, while the other parameters remain the same as the original ones after the approximations. Details on how to solve the RMF equations of motion to obtain the equation of state after these approximations are provided in our other work Guo et al. (2024b).